What is the volume of 0.1N HCl required to react completely with 1.0g of pure calcium carbonate

Grade:12th pass

1 Answers

Mahima Kanawat
6 years ago
Dear student ......for complete reaction the number of equivalents of the reactants must be same ...so equivalents of HCl = equivalents of CaCO3 ........Equivalents of HCl = N×V =0.1 × V ......Equivalent of CaCO3 = Wt / equivalent wt = 1/50 ( eq. Wt = mol wt./ valency factor = 100/2=50).....Hence , 0.1×V=0.02 .....therefore , V = 0.2 ltr or 200 ml.....With regards ...mahima...( Askiitians forum expert )